Secret:

Reduce the fraction to lowest terms. x^3 - x^2/x^3

Mercury:

are the x^3 and x^2 both part of the numerator? like this? \[\frac{ x^3-x^2 }{ x^3 }\]

Mercury:

or is it like this? \[x^3 - \frac{ x^2 }{ x^3 }\]

Secret:

No the top one, there both the numerator

Secret:

Their*

Mercury:

alright, cool you can split the fraction into x^3/x^3 - x^2/x^3 using the rule from before what is x^3/x^3? remember, top exponent minus bottom exponent

Secret:

it'll be 0 Right?

Mercury:

good so x^3/x^3 = x^0 which is equal to 1 using the same logic x^2/x^3 = ?

Secret:

x^1 Right?

Secret:

Wait no x^-1?

Mercury:

good, x^(-1) so x^3 - x^2/x^3 = 1 - x^(-1) or 1 - 1/x may I ask if you have a list of choices or do you type your answer in?

Secret:

I have to type in the answer..

Mercury:

hm. ok. let me just double check b/c they might want a certain format

Secret:

Oki

Mercury:

@hero may I ask your opinion on this question? we obtained the solution 1 - 1/x if the question prompt is "reduce to lowest terms" would it be better to write the answer as 1-1/x or (x-1)/x?

Hero:

Hint: Reducing to lowest terms implies writing as a single fraction.

Mercury:

so this? (x-1)/x?

Hero:

Correct

Mercury:

@secret can you try (x-1)/x as the solution? the parentheses are important, make sure to get them right

Secret:

Oki Thank you again!!! ^-^